OpenStudy (anonymous):

get a medal What is the x-coordinate of the solution of the following system of equations? 3x + y = 6 x - y = 6 -3 0 3

OpenStudy (imstuck):

In the second equation, y = x-6. Sub in that value for y in the first equation and get 3x+(x-6)=6. 4x-6=6. 4x=12. x=3. That's your answer.